Frederick C. Harris, Jr. Associate Professor

December 2, 2000

Deborah Poulson SIAM 3600 University City Science Center Philadelphia, PA 19104-2688

Dear Ms. Poulson,

Enclosed please find a copy of our paper entitled “Implementation of a Biologically Realistic Parallel Neocortical-Neural Network Simulator” which was accepted for publication in the Proceedings of the Tenth SIAM Conference on Parallel Processing for Scientific Computing.

We have tried to follow all the directions we were provided with, but if you find any problems with our paper or the ps file we will upload, please do not hesitate to call me or email me.

Thank you for your assistance in putting these proceedings together, and we look forward to possibly seeing you in March in Portsmouth Virginia.

Sincerely,

Frederick C. Harris, Jr.

fredh@cs.unr.edu
ph:  (775) 784-6571
fx:  (775) 784-1877